Grade 12Algebra

Set up equations and solve them to find the unknown numbers in the following cases:
Anwar thinks of a number. If he takes away 7 from 5/2 of the number, the result is 23.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us assume the required number be x
5/2 of the number = (5/2) x
The given above statement can be written in the equation form as,
= (5/2) x - 7 = 23
By transposing -7 from LHS to RHS it becomes 7
= (5/2) x = 23 + 7
= (5/2) x = 30 Multiply both side by 2,
= ((5/2) x) × 2 = 30 × 2
= 5x = 60
Then,
Divide both the side by 5 = 5x/5
= 60/5
= x = 12

Thanks